OpenStudy (anonymous):

Could someone show me how to simplify this. (n-4)!/(n-5)!

OpenStudy (anonymous):

@AkashdeepDeb @madrockz

OpenStudy (akashdeepdeb):

Good question! \[\frac{ (n-4)! }{ (n-5)! } \] = \[\frac{ (n-4)(n-5)(n-6)(n-7).........3.2.1 }{ (n-5)(n-6)(n-7).........3.2.1 }\] This is because factorials always reduce by 1! NOTE: n! = n(n-1)(n-2)(n-3)(n-4)...........3.2.1 So now cancel (n-5)! = \[\frac{ (n-4)(n-5)(n-6)(n-7).........3.2.1 }{~~~~~~~~~~~~~ (n-5)(n-6)(n-7).........3.2.1 }\] = \[\frac{(n-4)(n-5)!}{~~~~~~~~~~~~(n-5)!}\] It leaves you with \[(n-4)\] Understood? :)

OpenStudy (anonymous):

Okay, I would like help with one more in that step by step format you helped me in.... (n+4)!/(n+2)! and the problem (n+2)!/n!

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (akashdeepdeb):

If you understood this step by step format. Try the others yourself. I'll help.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

for (n+4)!/(n+2)! I got (n+2)(n+3)

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (anonymous):

Did I do it correctly? @AkashdeepDeb

OpenStudy (akashdeepdeb):

No see (n+4)! = (n+4)(n+3)(n+2)(n+1)n..............3.2.1 (n+2)! = (n+2)(n+1)n......3.2.1 Now divide them. What do you get?

OpenStudy (anonymous):

(n+2)?

OpenStudy (anonymous):

n-4

OpenStudy (akashdeepdeb):

No. \[\frac{ (n+4)(n+3)(n+2)(n+1)n..............3.2.1 }{ (n+2)(n+1)n......3.2.1 }\] DIVIDE! \[\frac{ (n+4)(n+3)(n+2)(n+1)n..............3.2.1 }{ ~~~~~~~~~~~~~~~~~~~~~(n+2)(n+1)n................3.2.1 }\] (n+2)! cancels out. Leaving us with \[(n+4)(n+3)\] Understood? :) Get the next one right! :COME ON! :D

OpenStudy (akashdeepdeb):

@BaileyC Did you **UNDERSTAND** this or not?

OpenStudy (anonymous):

Yes, i think i am solving the next question. please let me know when I am done solving it how i did

OpenStudy (anonymous):

i just told you

OpenStudy (akashdeepdeb):

Okay.

OpenStudy (anonymous):

(n+2)!/n! = (n+2)(n+1)

OpenStudy (akashdeepdeb):

Good! :)

OpenStudy (anonymous):

I really appreciate the help. R u a math teacher?

OpenStudy (akashdeepdeb):

Well! I am both. I learn while I teach. :) But not like a Pro Math teacher who earns! :P

OpenStudy (anonymous):

Whoa really? cool.

OpenStudy (anonymous):

Okay I have figured out the simplifying of expressions... Now I have a question on solving expressions. So I have the question here, n/(n-2)!=72 How do I solve for n with another number variable? @AkashdeepDeb

OpenStudy (akashdeepdeb):

Solve the LHS like you solved simply. And tell me the LHS.

OpenStudy (anonymous):

The answer is just 9 but I don't remember any of the steps to come to the single number! @AkashdeepDeb

OpenStudy (akashdeepdeb):

DO NOT MUG UP STEPS! Try to solve it by understanding factorials.

OpenStudy (anonymous):

hmm what do you mean

OpenStudy (akashdeepdeb):

http://www.mathsisfun.com/numbers/factorial.html

OpenStudy (anonymous):

I just dont get it algebraically

OpenStudy (akashdeepdeb):

n!/(n-2)! n(n-1)(n-2)!/(n-2)! [THIS IS THE FOURTH TIME] Plz SEE THE WEBSITE, and get your basics cleared. n(n-1) = 72 n(n-1) = 9*8 Thus n=9

OpenStudy (anonymous):

your mean

OpenStudy (akashdeepdeb):

*you're I speak the truth. See if you don't understand what I say, why don't you just ask me then? This is the 4th similar factorial problem. I am ready to help. Go through the website and I am sure you'll be able to solve these. :)

OpenStudy (anonymous):

I hope your video loses

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (akashdeepdeb):

OMG! Your curses aren't helping you. Neither are they moving me. :P Do you want to understand? :D

OpenStudy (anonymous):

bye